# Tropical Gradient Descent

### v2

@@ 3.2 Convergence Guarantees
$d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq\frac{2\alpha N}{\sqrt{m+1}-DNL(N-1)/4\alpha-1}.$
Proof We fix $m$ and define:
$M_{1}=\left\lfloor\xi(m)^{2}\right\rfloor.$
We note that:
$0\leq\sqrt{M_{1}}\leq\xi(m)$
So $\sqrt{M_{1}}$ lies between the roots of the polynomial
$p(t)=\frac{N}{L}t^{2}+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)t+\frac{N-LKN^{2}}{L}.$
@@
$\frac{N}{L}(M_{1}+1)+\left(\frac{KDN^{2}(N-1)}{4\alpha}-\frac{N\sqrt{m+1}}{L}\right)\sqrt{M_{1}}-KN^{2}\leq 0,$
$\frac{N\sqrt{M_{1}}\sqrt{M_{1}+1}}{L}+\frac{KDN^{2}(N-1)\sqrt{M_{1}}}{4\alpha}-\frac{N\sqrt{M_{1}}\sqrt{m+1}}{L}\leq KN^{2},$
$\frac{KDN(N-1)}{2}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
@@
For all $n\geq M_{1}$ and $m$ such that $1<\xi(m)$, we have $a_{n}\leq\epsilon_{m}/N$.
Proof We have assumed that $1<\xi(m)$, so:
@@
$\sqrt{M_{1}}$ $\geq\xi(m)-1,\geq(\sqrt{m+1}-KDNL(N-1)/4\alpha)-1.$
Hence, by the step size definition:
@@
$\leq\frac{2\alpha}{\sqrt{m+1}-KDNL(N-1)/4\alpha-1}=\epsilon_{m}/N.$
@@
For all $n\geq m$, we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
@@
$\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})$ $\leq\sum_{j}\Delta_{j}(\mathbf{t}_{M_{1}}),$
$=\sum_{k\leq K}\sum_{i\leq j\leq N}|t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}|,$
$\leq\sum_{k\leq K}\sum_{i\leq j\leq N}d_{{\rm tr}}(\mathbf{t}_{M_{1}},\mathbf{x}_{k}),$
$\leq\frac{KDN(N-1)}{2}.$
The partial sums are bounded by:
@@
$\leq-2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
@@
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq KDN(N-1)/2+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
$\leq\frac{2\alpha KN}{\sqrt{M_{1}}},$
$\leq K\epsilon_{m}.$
@@ 4 Numerical Experiments
In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and TrAdamax, with a focus on their stability behavior. The implementation for all these experiments is available in the GitHub repository at https://github.com/Roroast/TropicalGradDescent.
@@ 4.1 Methods
Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{n}$ for both methods, which we fix as
$a_{n}=\frac{\gamma\|\nabla f(\mathbf{t}_{n})\|}{\sqrt{n}},$
@@
Stochastic Gradient Descent. Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(t)$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{n}\in X$, $f(\mathbf{t};\mathbf{x}_{n})$. This is simplified with the notation $f_{n}(\mathbf{t})$. As with gradient descent, we use the step size sequence give by
$a_{n}=\frac{\gamma\|\nabla f_{n}(\mathbf{t}_{n})\|}{\sqrt{n}},$
@@
Adam Variants. The Adam algorithm (Kingma & Ba,, 2015) uses gradient momentum to compound consistent but small effects. This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights. The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
Algorithm 2 TrAdamax
$\mathbf{m},\mathbf{u}\leftarrow\mathbf{0}$;
for $n<$ max_steps do
     $\mathbf{g}_{ni}\leftarrow\nabla f_{i}(\mathbf{t}_{n-1})$;
     $\mathbf{d}_{ni}\leftarrow(\max_{j}(\mathbf{g}_{nj})-\min_{j}(\mathbf{g}_{nj}))*\mathbf{1}_{\mathbf{g}_{ni}>0}$;
     $\mathbf{m}_{n}\leftarrow\beta_{1}*\mathbf{m}_{n-1}+(1-\beta_{1})*\mathbf{d}_{n}$;
     $\mathbf{u}_{n}\leftarrow\max(\beta_{2}*\mathbf{u}_{n-1},|\mathbf{d}_{n}|)$;
     $\mathbf{t}_{n}\leftarrow\mathbf{t}_{n}-\alpha/(1-\beta_{1}^{n})*\mathbf{m}_{n}/\mathbf{u}_{n}$;
To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction. When looking to take higher moments of the gradient over $n$, the $L_{\infty}$ is the natural option for the tropical setting. We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{n}$:
$\mathbf{u}_{n+1}=\max(\beta_{2}\mathbf{u}_{n},\mathbf{d}_{n}).$
The tropicalized Adamax algorithm is then given by Algorithm 2.
4.2 Computation
We now look to solve our centrality statistics, Wasserstein projection problems, and tropical linear regression using the 7 gradient methods discussed: classical descent (CD), tropical descent (TD), stochastic gradient descent (SGD), tropical stochastic gradient descent(TSGD), Adam, Adamax, and TrAdamax. We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, a coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100. We take same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$. Appendix B provides further details on the simulated datasets and computational methodology used. For each gradient method, we use learning rates which have been optimized for each dataset size and dimension, across different distribution types (Appendix C).
Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima. We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, which can highlight certain error values at which several initializations have accumulated. When the CDF of one method dominates another, it has a higher probability of achieving the same error and is therefore a more suitable method.
As we will observe, tropical gradient methods are stable at fewer points than classical gradient methods; as a result, tropical methods typically have further to travel before reaching a stable point. Over few steps, this would be compensated for by selecting larger learning rates for tropical gradient methods relative to classical gradient methods, leading to less precise results in the case of tropical gradient methods. Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilise, given our initialisations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
4.2.1 Centrality Statistics
We first consider the behavior of tropical descent methods for centrality statistics, Fermat–Weber points and Fréchet means. These optimization problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods. The experimental behavior of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
Figure 5 shows the CDF of the log relative error over 1000 steps of each gradient method for the 12 datasets. We note that the deterministic methods achieve similar results for most datasets, while the stochastic methods achieve the worst error for all 12 datasets and TSGD is performing worse than SGD. The disparity between SGD and TSGD is greatest when the dimensionality of the data is greater than the size of the dataset, and for coalescent data. We see similar behavior in the log error of Tropical Descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy. While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (Ardila & Klivans,, 2006; Page et al.,, 2020) and may influence the uniqueness of minima.

### v3

@@ 3.2 Convergence Guarantees
$d_{{\rm tr}}(\mathbf{t}_{m},\mathbf{t}^{*})\leq{\frac{2\sqrt{2}\alpha N}{\sqrt{m+1}-DL/2\alpha+\sqrt{LN-1}-\sqrt{2}}=\epsilon^{\prime}_{m}.}$
In order to use Theorem 29, we will need to find some $M_{1}$ such that $a_{M_{1}}<\epsilon/N$ (our step size is small enough) and $s_{m}\geq\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-K\epsilon$ (we have taken enough steps to get close to $V$). Our condition that $m$ is large enough such that $\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}\geq\sqrt{2}$ ensures that such an $M_{1}$ exists, while $\sqrt{m}\geq 2\alpha N/D$ ensures that $M_{1}\leq m$. We can then take the largest such $M_{1}$ to ensure that the corresponding $\epsilon$ is as tight as possible.
Proof We fix $m$ satisfying 2 and define:
${\xi(m)}$ ${\coloneqq\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}\right)^{2}+LKN-1},}$
$M_{1}$ $\coloneqq\left\lfloor\xi(m)^{2}\right\rfloor.$
We first note that by the quadratic mean inequality and as $\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}\geq\sqrt{2}$:
$\xi(m)$ $=\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}\right)^{2}+LKN-1}$
$\geq\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\frac{1}{\sqrt{2}}\left(\frac{\sqrt{m+1}}{2}-\frac{KDL}{4\alpha}+\sqrt{LKN-1}\right)$
$\geq\frac{1}{\sqrt{2}}\left(\sqrt{m+1}-\frac{KDL}{2\alpha}+\sqrt{LKN-1}\right)\geq 1.$
Therefore we have that
${1}\leq\sqrt{M_{1}}\leq\xi(m),$
and so $\sqrt{M_{1}}$ lies between the roots of the polynomial
$p(t)=L^{-1}t^{2}+\left(KD/2\alpha-L^{-1}\sqrt{m+1}\right)t+L^{-1}-KN$}.
@@
${L^{-1}}(M_{1}+1)+\left({KD/2\alpha}-{L^{-1}}\sqrt{m+1}\right)\sqrt{M_{1}}-KN\leq 0,$
${L^{-1}}\sqrt{M_{1}}\sqrt{M_{1}+1}+{KD\sqrt{M_{1}}/2\alpha}-{L^{-1}}\sqrt{M_{1}}\sqrt{m+1}\leq KN,$
${KD}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1})\leq\frac{2\alpha KN}{\sqrt{M_{1}}}.$
We now show that $M_{1}\leq m$. Suppose otherwise; then by the inequality above:
$KD$ $\leq\frac{2\alpha KN}{\sqrt{M_{1}}}+2\alpha L^{-1}(\sqrt{m+1}-\sqrt{M_{1}+1})$
$<\frac{2\alpha KN}{\sqrt{m}}$
This contradicts our assumption that $m\geq(2\alpha N/D)^{2}$, so we conclude that $M_{1}\leq m$.
@@
For all $n\geq{m}$ we have $a_{n}\leq\epsilon_{m}/N$.
Proof We have shown that $\xi(m)\geq 1$, so:
@@
$\sqrt{M_{1}}$ $\geq\xi(m)-1,$
$\geq{\frac{1}{\sqrt{2}}\left(\sqrt{m+1}-\frac{KDL}{2\alpha}+\sqrt{LKN-1}\right)-1}.$
Hence, by the step size definition and the fact that $n\geq m\geq M_{1}$:
@@
$\leq{\frac{2\sqrt{2}\alpha}{\sqrt{m+1}-KDL/2\alpha+\sqrt{LKN-1}-\sqrt{2}}}=\epsilon_{m}/N.$
@@
We have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}\leq K\epsilon_{m}$.
@@
$\Delta_{j}(\mathbf{t}_{M_{1}})$ $=\sum_{k\leq K}\sum_{i\leq N}[t_{M_{1}j}-x_{kj}-t_{M_{1}i}+x_{ki}]^{+},$
$\leq\sum_{k\leq K}d_{\triangle_{\max}}(\mathbf{x}_{k},\mathbf{t}_{M_{1}}),$
$\leq KD.$
And therefore we have $\max_{j}\Delta_{j}(\mathbf{t}_{M_{1}})\leq KD$. The partial sums are bounded by:
@@
$\leq 2\alpha L^{-1}{(\sqrt{M_{1}+1}-\sqrt{m+1})}.$
@@
$\max_{i}\Delta_{i}(\mathbf{t}_{M_{1}})+s_{M_{1}}-s_{m}$ $\leq{KD}+2\alpha L^{-1}(\sqrt{M_{1}+1}-\sqrt{m+1}),$
$\leq\frac{2\alpha KN}{\sqrt{M_{1}}}\leq K\epsilon_{m}.$
@@
Our error bound in Proposition 12 is a useful foundational result on the convergence behavior of tropical descent; for sufficiently nice problems, tropical descent will converge at a rate of $O(1/\sqrt{m})$. However, a more detailed study of the numerical analysis of tropical descent would be valuable. We outline possible questions below, as well as presenting a comparison of our error bounds with tropical descent for tropical linear regression.
Firstly, the constant term $LD/2\alpha$ in the denominator of $\epsilon_{m}$ will typically be larger than necessary, resulting in a weaker error bound. In particular, this constant term is quick to dominate for larger problems. Further work improving these constant factors would produce much tighter and more robust error bounds. Secondly, the error $\epsilon_{m}$ bounds the distance to a tropical convex hull $V$, which is rarely a useful quantity to bound. We would rather use $\epsilon^{\prime}_{m}$, which bounds the distance to a global minimum, but our tropical location problems are not known to be $\triangle_{\min}$-star-quasi-convex. It would be particularly useful to generalize our error bound $\epsilon_{m}^{\prime}$ to such problems which exhibit some local (but not global) $\triangle_{\min}$-star-quasi-convexity.
For a qualitative view on Proposition 37, we compare $\epsilon_{m}^{\prime}$ and the convergence rate of tropical descent for tropical linear regression. As noted, tropical linear regression is tropically quasi-convex (Proposition 18) but not known to be $\triangle_{\min}$-star-quasi-convex, so does not meet the assumptions of Proposition 37; despite this, we see $\epsilon_{m}^{\prime}$ is often an effective upper bound on the error in $\mathbf{t}$.
Figure 5: The mean log error bound ${\rm log}\,\epsilon_{m}^{\prime}$ and the mean log tropical distance to the estimated minimum $\mathbf{t}^{*}$ when applying tropical descent (TD) with $\alpha=1$ to solve the tropical linear regression problem. For each dataset, we take 50 random initialization points and compute ${\rm log}\,\epsilon_{m}^{\prime}$ and ${\rm log}\,d_{\mathrm{tr}}(\mathbf{t}_{m},\mathbf{t}^{*})$ independently, then take an average. We also record the 10th and 90th percentile values to estimate an 80% confidence interval.
We perform tropical descent for tropical linear regression using 12 different datasets; these are generated via branching processes, coalescent processes, or a Gaussian distribution on $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$ and $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$. We consider sample sizes of 10 or 100. These are the same datasets used in Section 4, and further details on their generation can be found in Appendix B. We run computations using a learning rate of $\alpha=1$, for steps up to $m\leq 3000$. The tropical linear regression problem has $\|\nabla f(\mathbf{t}_{m})\|=2$ almost everywhere, so we take $L=1$ when computing $\epsilon_{m}^{\prime}$. We estimate $\mathbf{t}^{*}$ to be the mean of the final 10 steps of tropical descent (TD), and compute a tight estimate for $D$ by taking
$D\approx\max_{m\leq 3000}d_{\triangle_{\max}}(\mathbf{t}^{*},\mathbf{t}_{m}).$
Figure 5 shows the mean log error bound $\epsilon_{m}^{\prime}$ with the mean log tropical distance to $\mathbf{t}^{*}$.
Examining Figure 5, we first note that the error $\epsilon_{m}^{\prime}$ is orders of magnitude smaller for the lower dimensional problems, which is primarily a result of the constant terms in the denominator of $\epsilon_{m}^{\prime}$. As we have noted, it would be of significant value to reduce these constant terms in future numerical analytical studies of tropical descent. These constant terms are also the cause of noise in the mean error bound for high dimensional problems, as we require more steps for the error to be well-defined. We also note that although $\epsilon_{m}^{\prime}$ is a reasonable upper bound for the distance to $\mathbf{t}^{*}$ in many cases, this upper bound does not hold for the low dimensional coalescent datasets. It is unclear whither this is due to poor estimates for $D$ and $\mathbf{t}^{*}$, or due to tropical linear regression not being $\triangle_{\min}$-star-quasi-convex. We conclude by noting that we are observing a convergence rate of $O(1/\sqrt{m})$ for most experiments, with this rate appearing particularly distinctly for branching process and Gaussian data.
@@ 4 Numerical Experiments
In this section we compare the practical performance of classical descent, tropical descent, SGD, tropical SGD, Adam, Adamax, and tropical Adamax, with a focus on their stability behavior. The implementation for all these experiments is available in the GitHub repository at https://github.com/Roroast/TropicalGradDescent.
@@ 4.1 Methods
Steepest Descent. We will perform classical descent (CD) and tropical descent (TD) using equivalent step size sequences $a_{m}$ for both methods, which we fix as
$a_{m}=\frac{\gamma\|\nabla f(\mathbf{t}_{m})\|}{\sqrt{m}},$
where $\|\cdot\|$ is the Euclidean norm for classical descent and the tropical norm for tropical descent. When $\nabla f(\mathbf{t})$ is bounded this is a well-behaved step rule; $a_{m}$ will converge to 0. Furthermore, for the piecewise linear problems, $\|\nabla f(\mathbf{t}_{m})\|$ is bounded away from 0 almost everywhere which is not a global minimum, so $\sum a_{m}=\infty$.
Stochastic Gradient Descent. Stochastic gradient descent is a variant of gradient descent which looks to improve computational speed by approximating $\nabla f(\mathbf{t})$; rather than computing the loss and its derivative at $\mathbf{t}$ with respect to the entire dataset $f(\mathbf{t};X)$, we evaluate the loss and its derivative with respect to a random sample $\mathbf{x}_{m}\in X$, $f(\mathbf{t};\mathbf{x}_{m})$. This is simplified with the notation $f_{m}(\mathbf{t})$. As with gradient descent, we use the step size sequence given by
$a_{m}=\frac{\gamma\|\nabla f_{m}(\mathbf{t}_{m})\|}{\sqrt{m}},$
where $\|\cdot\|$ is the Euclidean norm for stochastic gradient descent (SGD) and the tropical norm for tropical stochastic gradient descent (TSGD).
Adam Variants. The Adam algorithm (kingma2017adam) uses gradient momentum to compound consistent but small effects. This is done by taking first and second moments of the gradient over all steps using exponentially decaying weights. The Adamax algorithm is a variant of the Adam algorithm motivated by the use of the $\infty$-moments rather than second moments.
To adapt the Adam algorithm for the tropical setting, we take moments of an unnormalized tropical descent direction. When looking to take higher moments of the gradient over $m$, the $L_{\infty}$ is the natural option for the tropical setting. We therefore look to use the Adamax recursion relation on the $\infty$-moment estimate $\mathbf{u}_{m}$:
$\mathbf{u}_{m+1}=\max(\beta_{2}\mathbf{u}_{m},\mathbf{d}_{m}).$
The tropicalized Adamax algorithm is then given by Algorithm 2.
Algorithm 2 TrAdamax
${\mathbf{v}},\mathbf{u}\leftarrow\mathbf{0}$;
for $m<$ max_steps do
  $\mathbf{g}_{mi}\leftarrow\nabla f(\mathbf{t}_{m-1})_{i}$;
  $\mathbf{d}_{mi}\leftarrow(\max_{j}(\mathbf{g}_{mj})-\min_{j}(\mathbf{g}_{mj}))*\mathbf{1}_{\mathbf{g}_{mi}>0}$;
  ${\mathbf{v}_{m}}\leftarrow\beta_{1}*{\mathbf{v}_{m-1}}+(1-\beta_{1})*\mathbf{d}_{m}$;
  $\mathbf{u}_{m}\leftarrow\max(\beta_{2}*\mathbf{u}_{m-1},|\mathbf{d}_{m}|)$;
  $\mathbf{t}_{m}\leftarrow\mathbf{t}_{m}-\alpha/(1-\beta_{1}^{m})*{\mathbf{v}_{m}}/\mathbf{u}_{m}$;
4.2 Computation
We now look to solve our centrality statistics, Wasserstein projection problems, and tropical linear regression using the 7 gradient methods discussed: classical descent (CD), tropical descent (TD), stochastic gradient descent (SGD), tropical stochastic gradient descent(TSGD), Adam, Adamax, and TrAdamax. We run computations for 12 datasets of different sizes, shapes and dimensions; motivated by phylogenetic data science applications, we consider datasets generated by branching processes, coalescent processes, or Gaussian distributions on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, for $N=6,28$ and samples of size 10 or 100. We take the same random initializations for each gradient method, taken from a Gaussian distribution on $\mathbb{R}^{N}$. Appendix B provides further details on the simulated datasets and computational methodology used. For each gradient method, we use learning rates which have been optimized for each dataset size and dimension, across different distribution types (Appendix C).
Rather than investigating the relative convergence rates of each method, we are interested in the relative stability of local minima. We therefore plot the CDF of the relative log error after 1000 steps over 50 random initializations, to reveal concentrations of error values where multiple initializations stabilize. When the CDF of one method dominates another, it has a higher probability of achieving the same error or less and is therefore a more suitable method.
As we will observe, tropical gradient methods are stable at fewer points than classical gradient methods; as a result, tropical methods typically have further to travel before reaching a stable point. Over few steps, this would be compensated for by selecting larger learning rates for tropical gradient methods relative to classical gradient methods, leading to less precise results in the case of tropical gradient methods. Our computations are run over 1000 steps, as this gives enough time for tropical gradient methods to stabilize, given our initializations, without sacrificing learning rate; we see in Appendix C that neither classical nor tropical learning rates are consistently greater than the other.
4.2.1 Centrality Statistics
We first consider the behavior of tropical descent methods for Fermat–Weber points and Fréchet means. These optimization problems are classically convex problems, as well as being both $\max$-tropical and $\min$-tropical location problems, and therefore lend themselves well to gradient methods. The experimental behavior of these problems is similar, so in this section we include the results for the Fermat–Weber problem while the results for Fréchet means can be found in Appendix E.
Figure 6 shows the CDF of the log relative error after 1000 steps of each gradient method for the 12 datasets. We note that the deterministic methods achieve similar results for most datasets, while the stochastic methods achieve the worst error for all 12 datasets and TSGD is performing worse than SGD. The disparity between SGD and TSGD is greatest when the dimensionality of the data is greater than the size of the dataset, and for coalescent data. We see similar behavior in the log error of tropical descent and TrAdamax; they perform on par with their classical counterparts for each dataset other than the 28 by 10 coalescent data, for which the tropical gradient methods are not achieving the same accuracy (Table 1). While it is unclear why this is, we note that coalescent data is supported on the space of ultrametric trees, which is a tropical hyperplane (ardila2006bergman; page2020tropical) and may influence the stability of minima.
Figure 6: The CDF of the log relative error after 1000 steps for each gradient method with tuned learning rates across 50 random (Gaussian) initializations when minimizing the Fermat–Weber objective. Each subfigure corresponds to a dataset with sample size $K=10,100$ of dimensionality $N=6,28$, sampled from a branching process, coalescent process, or a Gaussian distribution.
Table 1: The mean log relative error of each gradient method after 1000 steps with tuned learning rates for branching process, coalescent process, and Gaussian datasets of size $K=10,100$ in $\mathbb{R}^{6}/\mathbb{R}\mathbf{1}$, $\mathbb{R}^{28}/\mathbb{R}\mathbf{1}$ when minimizing the Fermat–Weber objective function. The minimal mean errors for each dataset are in bold.
Data CD TD SGD TSGD Adam Adamax TrAdamax
@@ 4.2.1 Centrality Statistics
[…]
4.2.2 Wasserstein Projections
The Wasserstein projection problems highlight the differences in stability behavior between classical and tropical gradient methods, as they are $\min$-tropical location problems but not classically convex. The CDFs of the final log relative errors can be found in Figures 7 and 8.